# BlendScape: Enabling Unified and Personalized Video-Conferencing Environments through Generative AI

### v1

@@ 3. Scenario Walkthrough
To illustrate how BlendScape can enable expressive meeting environments by blending physical and digital spaces, we present an initial scenario implemented with our system: a brainstorming session between two interaction designers who are collaborating to design a new mixed reality interface.
The designers join a video call and add their webcam feeds to BlendScape’s composition interface; one designer also adds a camera feed of their physical notebook in their design studio, so they can sketch ideas during the brainstorming session (Fig. 3a). The designers resize and reposition their videos through click-and-drag interactions, to render the whiteboard large and in the center of the environment. They type “brainstorming” in the prompt field to generate a creative environment for their planned design activity. A few seconds later, they see their blended environment: part of their physical surroundings are still visible, but they now extend into one unified design lab (Fig. 3b) where their own video feeds seamlessly blend in. This blended design allows them to gesture and refer to physical sketches in the notebook, simulating how they might interact in a co-located design studio. To preview mock-ups of their MR interface, the designers blend a digital sketch into their physical backgrounds (Fig. 3c).
Many other meeting experiences are possible (Appendix. A.2), from work scenarios to therapy sessions, birthday parties, vacation or wedding planning. As we describe in the next section, BlendScape provides flexible and modular techniques for creating such rich meeting spaces. Later, we discuss additional scenarios that demonstrate the expressiveness of BlendScape to enable distributed collaboration techniques from prior work (Sec. 5).
@@ 4. BlendScape System
We present BlendScape, a system for meeting participants to compose and customize video-conferencing environments. Key to our approach is blending elements of users’ physical or virtual backgrounds into unified environments, to incorporate visuals that are meaningful to the collaboration context or meeting participants. We first present three system requirements that guided the design of BlendScape. Then, we introduce a set of generative AI techniques for creating blended environments and composition techniques for enhancing the visual cohesion and naturally integrating users within the scene.
@@ 4.1. Requirements
We defined three requirements for generating blended environments with BlendScape, informed by our motivating scenario and our review of video-conferencing literature:
R1: Enabling embedding meeting context into the environment. Our Design Brainstorming scenario (Sec. 3) and review of environment composition techniques for distributed collaboration (Sec. 2.1) surfaced many examples where the environment is used to reflect the collaboration context, e.g., conveying a shared task space (Grønbæk et al., 2021) or establishing a meeting theme (Hunter et al., 2014). BlendScape enables this through an image inpainting technique, which generates new environments by blending users’ webcam backgrounds together, and an image-to-image technique, which transforms input images of overarching environments and composites users within it. We also implemented prompt enhancement techniques that utilize LLMs to encourage the generation of functional and stylistic elements to support specific meeting scenarios.
R2: Compositing users in a cohesive and immersive manner. To create the illusion of meeting in a shared space, we aimed to render users’ videos such that they appear naturally integrated with the environment. One key strategy is preserving a portion of users’ video backgrounds while generating blended spaces. This makes the region around the user match the physical conditions and noise present in their input video stream (i.e., maintaining realistic lighting, shadows, and color temperature). We also developed hidden surface removal techniques in BlendScape to place users behind foreground objects, obscuring the boundaries of their webcams.
R3: Enabling coarse- and fine-grained customization of environments. Prior video-conferencing tools demonstrated a need for meeting participants to update the entire scene (e.g., generating spatial layouts to support small vs. large group discussions (Grønbæk et al., 2023a)) as well as make granular adjustments (e.g., adding objects as visual aids (Hu et al., 2023b)). In addition to BlendScape’s coarse-grained inpainting and image-to-image techniques for generating new environments, we enable fine-grained adjustments by selecting portions of the scene to re-generate. We also maintain a history of generated environments to enable iteration on past results.
@@ 4.2. BlendScape Interface
Next, we provide an overview of the BlendScape system (Fig. 4). BlendScape’s user interface consists of a canvas in the center that displays blended environments and composites the users’ videos within them. Meeting participants can generate environments via two composition modes (Fig. 4a): leveraging the video feed from only their webcam (using inpainting), or refining the image that is already present in the canvas (using image-to-image techniques). This image on the canvas may be a result of a previous image generation step or an image uploaded by the user (Fig. 4f). Users can specify to what extent their video backgrounds should be preserved through a slider (Fig. 4i).
BlendScape allows users to enter two prompts for a Meeting Activity and Meeting Theme (Fig. 4c) to generate visuals relevant to the purpose of their meeting. To control whether the prompts or the users’ backgrounds should be prioritized in the environment generation, users can change the Prompt Strength (Fig. 4b). BlendScape offers direct manipulation techniques (via clicking, dragging, and pinching) as intuitive ways to position and scale users, both for steering the environment generation and placing users’ videos in the resulting scene.
To make fine-grained adjustments, users can add or remove objects by selecting a region of the environment (Fig. 4e) and specifying a prompt. BlendScape also offers automatic layout techniques (Fig. 4d) that composite users behind objects in the scene and automatically scale their videos to match the size of the objects. BlendScape saves a history of past environment generations and users’ positions within the scene, to enable iterating on past designs (Fig. 4g).
@@ 4.3. Generative AI Techniques for Blended Environments
BlendScape implements two classes of AI image generation techniques for composing blended video-conferencing environments: (1) an inpainting technique, which blends users’ physical or virtual backgrounds into a unified environment, and (2) an image-to-image technique, which incorporates users into an overarching environment that represents the meeting setting.
Generating blended environments with users’ video backgrounds (R1, R2)
With BlendScape’s inpainting techniques, meeting participants can blend their real-world surroundings or virtual backgrounds into the shared environment as a mechanism for personalization (Venolia et al., 2018) or capturing the task space where artifacts are collaboratively produced (Buxton, 2009). First, users can specify the proportion of their video backgrounds to retain; we mask these regions to preserve them in the blended environment (Fig. 6). Then, BlendScape performs inpainting to generate plausible visual details between the fixed regions, based on user-specified prompts of Meeting Activity and Meeting Theme, e.g., ‘‘brainstorming’’ in a ‘‘treehouse’’-themed environment (Fig. 5a). Preserving physical backgrounds, rather than generating fully artificial scenes, makes users appear more naturally integrated into the blended space, as their real-world conditions (e.g., lighting, shadows, color temperature, and webcam resolution) are reflected in the resulting scene.
To steer the generation of different styles of environments, users can reposition and scale their videos, e.g., positioning smaller images at the top of the screen can create a sense of depth, with a clear separation of foreground and background elements.
@@
Meeting participants can upload image priors to set the theme of the meeting, (e.g., using an image of a library for a study group meeting) or establish a layout of users (e.g., seating in rows for a college lecture). Furthermore, the image-to-image technique could be a strategy for re-creating a real-world meeting space that is familiar to distributed collaborators.
@@
To simplify novice users’ composition process into two clear pathways for creating environments, we separated the inpainting and image-to-image techniques into two modes within BlendScape: webcam-based and canvas-based generation. However, BlendScape supports flexibly combining these techniques to iterate on prior environments.
We found that performing an image-to-image transformation after an inpainting generation is an effective strategy for refining roughly-blended areas and subtly restyling an existing environment (Fig. 5c), through prototyping usage scenarios (Sec. 5) and observing participants’ design process in our User Study (Sec. 6). BlendScape also enables blending users’ video backgrounds into an image prior, to incorporate authentic elements of their backgrounds while adhering to an overarching environment structure.
@@ 4.4. Improving the Composition of Blended Environments
Next, we describe three composition techniques to improve the quality and make granular adjustments to the generated environments: (1) prompt-enhancement techniques to provide more detailed instructions to image generation models; (2) granular controls for scene-editing to add visuals or correct distortions; (3) hidden surface removal techniques to integrate users’ videos with foreground objects.
Generating context-informed and thematic environments through LLM-driven prompt padding (R2)
Crafting expressive prompts remains a challenge for novice users of image generation models, requiring significant trial and error (Brade et al., 2023; Chung and Adar, 2023b). To lower the barrier to entry for meeting participants to create high-quality and detailed environments, BlendScape enhances user-specified prompts with additional keywords suggested by large language models, a technique demonstrated in prior systems research (e.g., Opal (Liu et al., 2022)). We elicit text-based prompts from users in two steps to decouple requirements for the layout of the environment, which is driven by the Meeting Activity, and aesthetic elements that reflect the Meeting Theme. Then, query GPT-3.5 with the Meeting Activity and Meeting Theme prompts as input, instructing the model to return keywords for five relevant objects in the meeting space and five stylistic qualities that represent the meeting atmosphere. The advantage of this approach is dynamically eliciting keywords that are tailored to the meeting context while minimizing the effort and expertise required from meeting participants. For example, for a Meeting Activity of “Brainstorming Session” and Meeting Theme of “Hologram,” GPT-3.5 suggested “Interactive Touchscreens, Holographic Whiteboards Floating Desks as objects and “Dynamic Lighting Effects, Seamless Integration of Virtual and Physical Elements, Collaborative and Hi-tech Atmosphere” as stylistic qualities. Inspired by prompt expansion tools for artists (e.g., PromptGen555PromptGen: https://promptgen.vercel.app/), BlendScape also adds a fixed set of terms to encourage high-quality results: “highly detailed, intricate, sharp focus, smooth.” Appendix A.2 includes full examples of GPT queries and outputs.
@@
A limitation of capturing meeting participants via webcams is that they appear as “floating heads” with a harsh cut-off at their shoulders, due to the limited field-of-view. BlendScape takes a similar approach to address this issue as prior tools (e.g., Ohyay, TogetherMode): placing users behind objects in the scene and enabling hidden surface removal, a rendering technique that removes portions of 3D objects that should not be visible from a particular camera perspective. We implemented an object segmentation pipeline in BlendScape to compute the salient objects in the generated environment (e.g., chairs, tables) and extract them to a foreground layer positioned in front of users’ videos (Fig. 9). This creates the illusion of users sitting behind objects, hiding the bottom edge of their videos.
@@
In addition to BlendScape’s coarse-grained techniques for updating the entire video-conferencing environment, we implemented a selection technique for making fine-grained revisions to the scene (Fig. 7). Using the selection tool, users can circle areas of the scene and specify a text-based prompt to remove content (e.g., to fix distorted areas) or add visuals relevant to the meeting scenario (e.g., additional chairs to accommodate new meeting participants). To regenerate the area, BlendScape uses the GLIGEN (Li et al., 2023) inpainting model, which is trained to consider the prompt, position, and scale of the specified region in relation to its surroundings, to generate a more cohesive and realistic result.
@@ 4.5. Implementation
Image generation models: To enable the environment generation techniques, we used open-source Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) models through the WebUI API 666Stable Diffusion WebUI: https://github.com/AUTOMATIC1111/stable-diffusion-webui, ControlNet Extension: https://github.com/Mikubill/sd-webui-controlnet, hosted on a PC with an AMD EPYC 7742 64-Core Processor and NVIDIA RTX A6000 GPU. For the inpainting technique, we used the Realistic Vision 2.0 777https://huggingface.co/SG161222/Realistic_Vision_V2.0 Stable Diffusion checkpoint which is finetuned for inpainting, further guided by a ControlNet inpainting model. For the image-to-image technique, we used the Realistic Vision’s base checkpoint, guided by ControlNet Depth and Canny models to preserve the spatial layout and salient features of the image priors.
Furthermore, we incorporated GLIGEN (Li et al., 2023) to allow users to regenerate specific areas of the scene. Similar to ControlNet (Zhang et al., 2023b) in its goal of guiding diffusion models, GLIGEN makes region-specific edits based on textual prompts and bounding box coordinates.
Segmentation of the generated scene: BlendScape uses PixelLib (Olafenwa, 2021) to perform semantic segmentation to partition the scene into meaningful segments across the depth and for hidden surface removal. To segment users from their video backgrounds, we used conventional computer vision techniques. While BlendScape displays live video feeds of the segmented users, we generate static environments using the first frame of users’ videos, due to limitations with computing Stable Diffusion outputs at video rates. This results in an empty space in the video backgrounds when we perform person segmentation, since the users’ webcams do not capture areas that the user is occluding. We used Stable Diffusion to inpaint this area to simulate a continuous background.
2.5D, layered scenes: The inpainting and image-to-image techniques require isolating different components of the scene (i.e., image-to-image transforms the entire existing environment while inpainting only takes the users’ webcam backgrounds as input). To assemble the correct input for the image generation models, we implemented BlendScape as a 2.5D scene in Unity with five 2D layers staggered at different depths: (1) foreground objects, (2) users’ videos (isolated from their video backgrounds), (3) users’ video backgrounds, (4) background masks (i.e., to preserve regions of the video backgrounds), and (5) the environment (as shown in Fig. 9). We instrument the Unity scene with multiple orthographic cameras that only render specific layers, enabling us to capture each layer separately. Since orthographic cameras do not perform perspective rendering, scene elements are rendered at a consistent scale even when placed at different depths.

### v2

@@ 3. Scenario Walkthrough
To illustrate how BlendScape can enable expressive meeting environments by blending physical and digital spaces, we present a scenario implemented with our system: a brainstorming session between two designers who are prototyping a mixed reality interface.
The designers join a video call and add their webcam feeds to BlendScape’s composition interface. One designer also adds a camera feed of their physical notebook, so they can sketch ideas during the brainstorming session (Fig. 3a). At the start of the meeting, they use BlendScape to generate a creative environment for their ideation activity. They type “brainstorming” in the prompt field, and a few seconds later, they see their blended environment: their physical surroundings are still visible, but now seamlessly extend into a unified design studio (Fig. 3b). This blended meeting space allows them to gesture and refer to physical sketches in the notebook, simulating how they might collaborate face-to-face. Later in the meeting, the designers blend a digital sketch into their physical backgrounds to preview mock-ups of their MR interface (Fig. 3c).
Many other meeting experiences are possible, from work scenarios to therapy spaces, birthday parties, or vacation planning (Appendix. A.2). As we describe next, BlendScape provides flexible techniques for creating rich meeting spaces. Later, we present additional scenarios that demonstrate BlendScape’s expressiveness to enable distributed collaboration techniques from prior work (Sec. 5).
@@ 4. BlendScape System
This section presents BlendScape, a rendering and composition system that enables meeting participants to customize video-conferencing environments. Key to our approach is blending elements of users’ physical or virtual backgrounds into unified environments to allow for flexible representations of task spaces. We first outline three system requirements that guided the design of our system. Then, we describe our implementation of BlendScape’s generative AI techniques for customizing meeting environments and composition techniques for enhancing visual cohesion.
@@ 4.1. Requirements
We defined three requirements for customizing meeting environments with BlendScape, informed by our review of environment composition techniques for distributed collaboration (Sec. 2.1) and our motivating scenarios (Sec. 3, 5):
R1: Enabling users to express the meeting context through the environment. Embedding relevant visual and structural elements within video-conferencing environments can strengthen distributed users’ collaborative processes (e.g., conveying shared task spaces in our Design Brainstorming scenario, using spatial landmarks to facilitate conversational transitions (Hu et al., 2023a; Tang et al., 2023)). BlendScape allows users to align environments to their meeting purpose by specifying text-based prompts and providing image priors (i.e., images of their physical surroundings or other spaces that represent their collaboration needs).
R2: Supporting convincing illusions of meeting in a shared space. To simulate face-to-face communication cues (e.g., deictic gestures (Grønbæk et al., 2021), making eye-contact (Tang et al., 2023)), prior systems composite users within virtual environments that resemble physical spaces. To enable these designs while allowing users to incorporate their physical context, BlendScape implements two rendering techniques: (1) preserving and blending users’ video backgrounds into a unified environment, which maintains realistic lighting, color temperature, and shadows around users; (2) hidden surface removal to obscure the boundaries of users’ webcams among objects in the scene.
R3: Enabling coarse- and fine-grained customization of environments. Adapting meeting spaces to dynamic collaboration needs may require users to update the entire scene (e.g., to reconfigure the space for small vs. large group discussions (Grønbæk et al., 2023a)), as well as make minor adjustments (e.g., incorporating shared content (Hu et al., 2023b)). In addition to inpainting and image-to-image techniques for creating new environments, BlendScape enables users to make granular changes by selecting portions of the scene to re-generate. To enable iteration on past results, we maintain a history of environments.
@@ 4.2. BlendScape Interface
Next, we provide an overview of BlendScape’s composition tools (Fig. 4). BlendScape’s user interface consists of a canvas in the center that displays blended environments and composites the users’ videos within them. All meeting participants share the same view of this canvas; changes to the environment and position of users’ video feeds are synchronized across all users.
Meeting participants can generate environments using two composition modes (Fig. 4a): blending the video feeds from only their webcams (using inpainting), or refining the image that is already present in the canvas (using image-to-image techniques). This image on the canvas may be a result of a previous image generation step or an image uploaded by the user (Fig. 4f). For inpainting, users can specify how much of their video backgrounds should be preserved through a slider (Fig. 4i).
To generate visuals relevant to the purpose of their meeting, users can enter two prompts for a Meeting Activity and Meeting Theme (Fig. 4c). The Prompt Strength slider controls the prompt weight, i.e., how much to prioritize the prompts in the environment generation (Fig. 4b). BlendScape offers direct manipulation techniques (clicking, dragging, and pinching) as intuitive ways to position and scale videos, both for steering the environment generation and placing users’ videos in the scene.
To make fine-grained adjustments to the environment, users can add or remove objects by selecting a region of the scene (Fig. 4e) and specifying a prompt. BlendScape also offers automatic layout techniques (Fig. 4d) that composite users behind objects in the scene and automatically scale their videos to match the size of the objects. BlendScape saves a history of past environment generations and users’ positions within the scene, to enable iterating on past designs (Fig. 4g).
@@ 4.3. Generative AI Techniques for Blended Environments
To enable real-time end-user customization of video-conferencing environments, BlendScape implements two classes of AI image generation techniques: (1) an inpainting technique, which blends users’ physical or virtual backgrounds into a unified environment, and (2) an image-to-image technique, which incorporates users into an overarching image that represents the meeting setting.
Generating blended environments from users’ video backgrounds (R1, R2)
BlendScape allows meeting participants to incorporate their real-world surroundings or virtual backgrounds into the shared environment as a mechanism for personalization (Venolia et al., 2018) or capturing the task space where artifacts are collaboratively produced (Buxton, 2009). First, users can specify the proportion of their video backgrounds to retain; we mask these regions to preserve them in the blended environment (Fig. 6). Then, BlendScape performs inpainting to generate plausible visual details between the fixed regions, based on user-specified prompts for the Meeting Activity and Meeting Theme, e.g., ‘‘brainstorming’’ in a ‘‘treehouse’’-themed environment (Fig. 5a). By preserving physical backgrounds, we aimed to more naturally integrate users into the blended space by matching their real-world conditions (e.g., lighting, shadows, color temperature, and webcam resolution).
Direct manipulation of video feeds (i.e., re-positioning and re-scaling) can be used to steer the generation of different styles of inpainted environments. For example, positioning smaller videos at the top of the screen creates a sense of environmental depth, with clear separation of foreground and background elements.
@@
Meeting participants can upload image priors to set the theme of the meeting, (e.g., using an image of a library for a study group), establish a layout of users (e.g., row-based seating for academic lectures), or situate themselves in real-world meeting spaces.
@@
To simplify novice users’ composition processes into two clear pathways for creating environments, we separated the inpainting and image-to-image techniques into two modes within BlendScape: webcam-based and canvas-based generation. However, BlendScape supports flexibly combining these techniques to iterate on prior environments. For example, performing an image-to-image transformation after an inpainting generation is an effective strategy for refining roughly-blended areas and subtly restyling an existing environment.
@@ 4.4. Improving the Composition of Blended Environments
To make the base environments generated via inpainting and image-to-image techniques more suitable for video-conferencing, we implemented three composition techniques: (1) prompt enhancement to further align environments to the meeting context and enhance the visual quality; (2) granular scene-editing controls to add visuals or correct distortions; (3) hidden surface removal techniques to integrate users’ videos with foreground objects.
Generating context-informed and thematic environments through LLM-driven prompt padding (R1, R2)
Crafting expressive prompts remains a challenge for novice users of image generation models, requiring significant trial and error (Brade et al., 2023; Chung and Adar, 2023b). To lower the barrier for meeting participants to create detailed environments, BlendScape enhances user-specified prompts with keywords suggested by LLMs, as demonstrated by prior systems (Liu et al., 2022; Brade et al., 2023).
First, we elicit text-based prompts from users to define the Meeting Activity, which drives the layout of the environment, and the Meeting Theme, which steers the generation of aesthetic elements. Then, to dynamically tailor the image generation prompts to the meeting context, we query GPT-3.5 to augment the Meeting Activity and Meeting Theme prompts with keywords for five relevant objects and five stylistic qualities that represent the meeting atmosphere. For example, for a Meeting Activity of “Brainstorming Session” and Meeting Theme of “Hologram,” GPT-3.5 suggested “Interactive Touchscreens” and “Holographic Whiteboards” as objects and “Dynamic Lighting” and “Seamless Integration of Virtual & Physical Elements” as stylistic qualities.
Inspired by prompt expansion tools for artists (e.g., PromptGen333PromptGen: https://promptgen.vercel.app/), BlendScape also adds a fixed set of terms to encourage high-quality visuals: “highly detailed, intricate, sharp focus, smooth.” Appendix A.2 includes examples of GPT queries and outputs.
@@
A limitation of capturing meeting participants via webcams is that they appear as “floating heads” with a harsh cut-off at their shoulders, due to the limited field-of-view. To more naturally integrate users within environments, BlendScape takes a similar approach as prior tools (e.g., Ohyay, TogetherMode11footnotemark: 1): placing users behind objects in the scene and enabling hidden surface removal, a rendering technique to remove portions of 3D objects that should not be visible from a particular camera perspective. We implemented an object segmentation pipeline in BlendScape to compute the salient objects in the generated environment (e.g., chairs, tables) and extract them to a foreground layer positioned in front of users’ videos (Fig. 9). This creates the illusion of users sitting behind objects.
@@
In addition to BlendScape’s inpainting and image-to-image for updating the entire video-conferencing environment, we implemented a technique for making fine-grained revisions to the scene (Fig. 7). Using the selection tool, users can circle areas of the scene and specify a text-based prompt to remove content (e.g., to fix distorted areas) or add visuals relevant to the meeting scenario (e.g., chairs to accommodate new meeting participants). To regenerate the area, BlendScape uses the GLIGEN (Li et al., 2023) inpainting model, which is trained to generate cohesive results by considering the prompt, position, and scale of the specified region in relation to its surroundings.
@@ 4.5. Implementation
Image generation models: To enable the environment generation techniques, we used open-source Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) models through the WebUI API444Stable Diffusion WebUI: https://github.com/AUTOMATIC1111/stable-diffusion-webui, ControlNet Extension: https://github.com/Mikubill/sd-webui-controlnet, hosted on a PC with an AMD EPYC 7742 64-Core Processor and NVIDIA RTX A6000 GPU. For inpainting, we used the Realistic Vision 2.0555Realistic Vision 2.0: https://huggingface.co/SG161222/Realistic_Vision_V2.0 Stable Diffusion checkpoint which is fine-tuned for inpainting, further guided by a ControlNet inpainting model. For image-to-image generations, we used Realistic Vision’s base checkpoint, guided by ControlNet Depth and Canny models to preserve the spatial layout and salient features of the image priors. Furthermore, we incorporated GLIGEN (Li et al., 2023) to allow users to regenerate specific areas of the scene. Similar to ControlNet (Zhang et al., 2023b) in its goal of guiding diffusion models, GLIGEN makes region-specific edits based on textual prompts and bounding box coordinates.
Using these models, the average generation times were 10s for inpainted environments, 25s for image-to-image environments, and 20s for GLIGEN-enabled edits. This includes the time to enhance users’ prompts with additional keywords (via GPT-3.5).
2.5D, layered scenes: The image-to-image generation mode transforms the entire existing environment, while the inpainting mode only takes users’ webcam backgrounds as input. To isolate the correct scene elements for the image generation models, we implemented BlendScape as a 2.5D scene in Unity with five 2D layers staggered at different depths (Fig. 9): (1) foreground objects, (2) users’ videos (separated from their video backgrounds), (3) users’ video backgrounds, (4) background masks (to preserve regions of the video backgrounds), and (5) the generated environment. We instrumented the Unity scene with multiple cameras that render specific layers, enabling us to capture each layer separately. We use orthographic cameras (i.e., cameras that do not perform perspective rendering), so scene elements are rendered at a consistent scale even when placed at different depths.
Environment segmentation: BlendScape uses the PixelLib (Olafenwa, 2021) semantic segmentation model to partition scene elements for hidden surface removal. We segment users from their video backgrounds via conventional computer vision techniques. While BlendScape displays live video feeds of the segmented users, we generate static environments using the first frame of users’ videos, due to limitations with running Stable Diffusion at video rates.
Performing person segmentation leaves empty spaces in static video backgrounds, as webcams fail to capture areas of the background that are occluded by the user. We used Stable Diffusion to inpaint this area, simulating a continuous background (Fig. 8).